OpenStudy (anonymous):

Determine whether the following series converges or diverges. (-1)^(n-1)/sqrt(n)

OpenStudy (anonymous):

alternating, so as long as the terms go to zero it converges

OpenStudy (anonymous):

i believe this converges because of the alternating series but i wanted to ask what do you have to look at to know if it converges when there is a square root in the denominator?

OpenStudy (anonymous):

by "series" i assume you mean \[\sum_{k=1}^{\infty}\frac{(-1)^{k+1}}{\sqrt{k}}\]

OpenStudy (anonymous):

converges because a) it is alternating and b) \[\lim_{n\rightarrow \infty}\frac{1}{\sqrt{n}}=0\]

OpenStudy (anonymous):

yes that is what i mean

OpenStudy (anonymous):

ohh so as the nth term gets bigger in the denominator the limit gets closer and closer to zero, correct?